Steve traveled 150 miles at a certain speed. Had he gone 20mph faster, the trip would have taken 2 hours less. Find the speed of his vehicle. Can you help me with this one?
Answer by edjones(8007)

st=d
st=150
(s+20)(t-2)=150
.
t=150/s
(s+20)(150/s -2)=150
(s+20)((150-2s)/s)=150
(s+20)(150-2s)=150s
-(s+20)(2s-150)=150s
-(2s^2-110s-3000)=150s
-2s^2-40s+3000=0
s^2+20s-1500=0
(s+50)(s-30)=0
s=30
.
Ed